# Security and Privacy on Generative Data in AIGC: A Survey

### v2

@@ IV-B2 Blockchain
Du et al. [72] proposed a blockchain-empowered framework to manage the lifecycle of AIGC-generated data. Firstly, a protocol to protect the ownership and copyright of AIGC is proposed, called Proof-of-AIGC, deregistering plagiarized generative data and protecting users’ copyright. Then, they designed an incentive mechanism with one-way incentives and two-way guarantees to ensure the legal and timely execution of AIGC ownership exchange funds between anonymous users.
IV-C Summary
In table III, we summarize the solutions for the controllability of generative data in AIGC. We emphasize the discussion on the access control and traceability of generative data. By implementing them, we can protect the security and privacy of generative data, ensuring its credibility and reliability and providing robust support for the compliant use of the data. Adding adversarial perturbations and watermarks to data both involve the introduction of additional information. However, these additional pieces of information are susceptible to removal by existing adversarial techniques [73]. Therefore, it is crucial to enhance the robustness of these technologies, which will increase their effectiveness against all forms of attacks, while also ensuring data integrity, authenticity and persistence of ownership.
V Authenticity on Generative Data
V-A Threats to Authenticity
@@
Dramatic advances in generative models have made significant progress in generating realistic data, reducing the amount of expertise and effort required to generate fake content. However, this unrestricted accessibility raises concerns about the ubiquitous spread of misinformation. Fake images are particularly convincing due to their visual comprehensibility. As a result, malicious users can generate harmful content to manipulate public opinion, thereby negatively impacting social domains, e.g., politics and economics. For example, the fake tweet with the generated image of “a large explosion near the Pentagon complex” went viral, fooling many authoritative media accounts into reprinting them and even causing the stock market to suffer a significant drop. Current state-of-the-art generative models already pose a greater threat to human visual perception and discrimination. When distinguishing between real images and generative images, the error rate of human observers reaches 38.7% [74]. Typically, Deepfake [75] possesses the capacity to fabricate visually realistic fake content by grafting the identity of one individual onto the image or video of another. Unfortunately, the open-source nature of the technology allows criminals to commit malicious forgeries without the need for significant expertise, thereby engendering a multitude of societal risks [11, 76]. For instance, this includes replacing the protagonist in pornographic videos with a celebrity face to affect the celebrity’s reputation, faking videos of speeches of politicians to manipulate national politics, and faking an individual’s facial features to pass authentication in assets management. Many fake detection methods [77] have been proposed to detect modified data by AI. However, these methods still have vulnerabilities and limitations [78]. On the one hand, some methods rely too heavily on traditional principles or pattern matching. They are difficult to capture the evolving new patterns of in-depth AIGC, which allows generative data to escape detection. On the other hand, existing methods have limited capabilities when dealing with new challenges under large models. Large models have higher generative power and creativity, making the generative data more difficult to distinguish. A recent study [79] provided insights into the various methods used to detect ChatGPT-generated text. The study highlighted the extraordinary ability of Chat-GPT spoofing detectors and further shows that most of the analyzed detectors tend to classify any text as human-written with an overall TNR as high as 90% and a low TPR. Therefore, there is a requirement to continuously improve the existing detector to effectively deal with the problem of disinformation and misuse of generative data.
V-B Countermeasures to Authenticity
@@
In Fig. 7, existing countermeasures mainly consider constructing a detector to distinguish between real data and generative data. Further, generative attribution can trace the generative data back to the model that generate it.
Figure 7: Detection and attribution of generative data.
V-B1 Generative Detection

### v3

@@ 4.2.2. Blockchain
Du et al. (Liu et al., 2024b) proposed a blockchain-empowered framework to manage the lifecycle of AIGC-generated data. Firstly, a protocol to protect the ownership and copyright of AIGC is proposed, called Proof-of-AIGC, deregistering plagiarized generative data and protecting users’ copyright. Then, they designed an incentive mechanism with one-way incentives and two-way guarantees to ensure the legal and timely execution of AIGC ownership exchange funds between anonymous users. AIGC-Chain(Jiang et al., 2024) carefully records the entire life cycle of AIGC products, providing a transparent and reliable platform for copyright management.
4.3. Summary
In table 3, we summarize the solutions for the controllability of generative data in AIGC. We emphasize the discussion on the access control and traceability of generative data. By implementing them, we can protect the security and privacy of generative data, ensuring its credibility and reliability and providing robust support for the compliant use of the data. In addition, adversarial attacks can also have an impact on the controllability of generative data. On the one hand, adversarial attacks can attack the watermark remover to prevent the loss of controllable information, which achieves stable controllability of generative data. On the other hand, adversarial attacks can also attack the watermark extractor to prevent the extraction of controllable information, which exacerbates the controllability challenge on generative data.
5. Authenticity on Generative Data
5.1. Threats to Authenticity
@@
Dramatic advances in generative models have made significant progress in generating realistic data, reducing the amount of expertise and effort required to generate fake content. However, this unrestricted accessibility raises concerns about the ubiquitous spread of misinformation. Fake images are particularly convincing due to their visual comprehensibility. As a result, malicious users can generate harmful content to manipulate public opinion, thereby negatively impacting social domains, e.g., politics and economics. For example, the fake tweet with the generated image of “a large explosion near the Pentagon complex” went viral, fooling many authoritative media accounts into reprinting them and even causing the stock market to suffer a significant drop. Current state-of-the-art generative models already pose a greater threat to human visual perception and discrimination. When distinguishing between real images and generative images, the error rate of human observers reaches 38.7% (Lu et al., 2023a).
Typically, Deepfake (Mirsky and Lee, 2021) possesses the capacity to fabricate visually realistic fake content by grafting the identity of one individual onto the image or video of another. Unfortunately, the open-source nature of the technology allows criminals to commit malicious forgeries without the need for significant expertise, thereby engendering a multitude of societal risks (Korshunov and Marcel, 2018; Aliman and Kester, 2022). For instance, this includes replacing the protagonist in pornographic videos with a celebrity face to affect the celebrity’s reputation, faking videos of speeches of politicians to manipulate national politics, and faking an individual’s facial features to pass authentication in assets management.
Many fake detection methods (Verdoliva, 2020) have been proposed to detect modified data by AI. However, these methods still have vulnerabilities and limitations (Hussain et al., 2021). On the one hand, some methods rely too heavily on traditional principles or pattern matching. They are difficult to capture the evolving new patterns of in-depth AIGC, which allows generative data to escape detection. On the other hand, existing methods have limited capabilities when dealing with new challenges under large models. Large models have higher generative power and creativity, making the generative data more difficult to distinguish. A recent study (Pegoraro et al., 2023) provided insights into the various methods used to detect ChatGPT-generated text. The study highlighted the extraordinary ability of Chat-GPT spoofing detectors and further shows that most of the analyzed detectors tend to classify any text as human-written with an overall TNR as high as 90% and a low TPR. Therefore, there is a requirement to continuously improve the existing detector to effectively deal with the problem of disinformation and misuse of generative data.
5.2. Countermeasures to Authenticity
@@
In Fig. 7, existing countermeasures (Lin et al., 2024) mainly consider constructing a detector to distinguish between real data and generative data. Further, generative attribution can trace the generative data back to the model that generate it.
Figure 7. Detection and attribution of generative data.
5.2.1. Generative Detection